# Extended thermal cycling of ATLAS ITk strip modules with and without stress mitigating interposers

### v1

@@ 1 Introduction
The successful upgrade of the ATLAS Experiment [1, 2] for the High-Luminosity Large Hadron Collider (HL-LHC) [3] underpins future discoveries in particle physics [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Its cornerstone is a new all-silicon Inner Tracker (ITk) detector [16, 17, 18, 19]. The ITk comprises pixel [17] and strip [18] sub-detectors, each designed as a barrel bookended by endcaps. The strips barrel requires detector units called modules, each comprising mm p-type silicon strip sensors fabricated by Hamamatsu Photonics K. K. (HPK) [20]. Industrial adhesives [21, 22, 23] mechanically attach sensors to custom electronics supplying front-end power, readout, monitoring and control [24, 25, 26, 27, 28, 29]. Module assembly and testing within operating expectations, called quality control (QC) [30, 31], are distributed worldwide.
An important production QC procedure is module thermal cycling (MTC). Modules must pass electrical tests during repeated cycles between two temperatures. The nominal operating window is between room (C) and end-of-life temperatures (C) to mitigate elevated leakage currents from radiation damage [32, 33]. Every module undergoes ten thermocycles to verify electrical functionality and mechanical integrity between C. Cooling or power failures [34] can expose modules to temperatures outside the operating window. This motivates studies beyond nominal QC temperatures. Testing outside operating expectations is called quality assurance (QA).
However, we observe critical problems when thermocycling ten pre-production modules between C and C: visual inspection identifies a sensor fracture for one module (figure 1) and three modules develop excess leakage currents during C tests. Independent efforts report similar failure modes but only after loading modules onto local support structures, which is called the sensor fracturing problem [35, 36]. Simulation studies [37] suggest adverse stress from mismatched coefficients of thermal expansion (CTE) [38, 39, 40, 41] between the silicon sensor (fixed to local supports with adhesives) and copper in the flex printed circuit board (PCB); the intermediate epoxy glue provides a high-modulus mechanical coupling that transfers stress.
@@ 2 Module assembly and setup
We briefly review the pre-production module assembly procedure and thermocycling setup. Figure 3(a) illustrates the original pre-production barrel module design. Pre-production assembly uses Loctite Eccobond F112 [23] epoxy resin111The ITk Strips community refers to this adhesive as TrueBlue due to its optical colour.. This attaches the sensor to hybrid(s) and powerboard PCBs housing custom microelectronics [53, 54, 55, 56] for front-end readout, control, and power. There are two types of modules: long strip (LS) that have one hybrid to read out 2560 channels, and short strip (SS) that have two hybrids to read out 5120 channels. After wire bonding, modules undergo QC testing [30] in the coldbox of the MTC setup situated at the University of Cambridge (figure 4). This apparatus combines thermoelectric Peltiers with liquid coolant circulated by a Grant chiller [57, 58, 59], interfaced with custom detector control system (DCS) and data acquisition (DAQ) [60, 61], with continuous environmental monitoring.
This paper probes mechanical deformations by measuring the sensor bow of free modules outside the MTC setup. We employ the Baty Venture 3030 coordinate-measuring machine (CMM) [62]. We develop an in-house script to convert the machine output into a Common Data Format (CDF) defined by standard QC procedures [30, 43]. A community-developed script then reads this CDF file and computes the sensor bow using a standard bow algorithm. The algorithm first performs a fit to all sensor points to define a sensor plane. The heights of all sensor points are defined relative to this plane. The bow is defined as
@@ 3.2 High-voltage supply and interlock
We also develop a custom HV power supply as an add on to the MTC setup. It is based on a Spellman MPS DC-to-HV converter capable of delivering 2.5 mA at V. It uses an Arduino Leonardo microprocessor to operate the display, facilitate over-current and safety interlock functionality, and provide a serial-over-USB interface for remote readout and control. The HV converter has monitoring outputs for voltage and current, which are accurate to 1% of the full scale only. The unit features a dedicated interlock by interfacing to a magnetic reed switch attached to the coldbox lid (figure 5). When the lid is opened with HV enabled, the interlock rapidly reduces the HV output to 0 V. This introduces added safety beyond the existing software interlock.
@@ 3.3 Improved thermal chuck flatness
After fitting the shims and electrical connection pieces between the shims, feeler gauges determine the residual height differences with results in table 1. The post-shim values are all lower relative to the corresponding point before shimming, where the mean reduces to m with a maximum 170 m. This demonstrates improved chuck flatness relative to the original design. A bow of up to m is deemed tolerable for subsequent use. Comparative metrology measurements indicate a high degree of similarity between the bow of a module resting on a flat gauge plate and a module resting on a test frame bolted to a known flat jig surface. For consistency in the upcoming studies, sensor bow metrology is performed with the module resting on the same test frame.
![[Uncaptioned image]](x6.png)
@@ 4.1 Test-to-destruction methodology
Extend thermocycling until fractures form. Sensor deformation from thermal stress accumulates in the sensor bow with repeated MTC cycles, which can lead to sensor fractures. We evaluate electrical data and perform sensor visual inspection to search for fracture formation.
@@ 4.2 Bow evolution versus cycling temperature
We first report the regime with no sensor deformation. After the first cycle between C, we find a minor bow decrease of around m and the trends are correlated across all five modules, suggesting mechanical relaxation. For simplicity, we average the bow data over the five modules after C MTC to yield m (table 3). Thermocycling between C induces little change in bow, except LS-06 rises anomalously to m after the second C cycle due to accidental sensor damage. Leaving the modules in dry storage for five days decreases the average bow by m, indicating mechanical relaxation. After C thermocycling, we find the mean bow reaches m, where the average of the individual increases is m. The large standard deviation arises from the anomalous LS-06 value; excluding this outlier yields m. This establishes no statistically significant change in sensor deformation during MTC from C to C.
@@
This demonstrates that thermocycling above a critical temperature C “bakes in” sensor deformations permanently, enabling ex situ measurement outside the MTC setup. This hysteresis is also reported in endcap modules [43] and is hypothesised to arise from a glass transition of the epoxy adhesive [44] at C. The mechanism of how the glass transition induces sensor deformation is not fully understood and remains an empirical observation for this paper.
@@ 4.3 Post-mortem of sensor fractures
Differences in fracture morphology. The LS-02 sensor distinctly fractures into multiple pieces, where we visually identify five major segments. We refer to the fractures that do not propagate from one edge to the opposite edge as secondary fractures. We observe that these develop perpendicular to the primary fracture left of the powerboard shield box. One secondary fracture (medium blue) follows the powerboard glue edge with three near-right-angle kinks turning to the sensor top edge. The secondary fracture (dark blue) arises just above the primary fracture by the hybrid before turning perpendicularly at the top-left corner of the powerboard shield box to the sensor top edge, which does not trace a glue boundary.
@@ 5 Sensor deformation with stress-mitigating interposer
Simulation studies [37] suggest an order-of-magnitude reduction in stress by adding Kapton-silicone interposers. Due to its low modulus, the silicone gel reduces mechanical stress by decoupling the thermal deformation of the flexes from that of the sensor during cooling [46, 47]. Figure 10(a) schematically shows the flex-epoxy-sensor stackup augmented with 50 m of Kapton film and 100 m of SE-4445 silicone gel, whose concept is introduced in Refs. [37, 48]. These two materials are already validated for use in the ITk detector because SE-4445 attaches modules onto local supports, whose electrical services supplied by bus tapes are made of Kapton [65]. Introducing these extra thin layers across all modules has a negligible impact on ITk material budget. This motivates us to prototype interposers in our laboratory by developing in-house designs, ad hoc assembly procedures, tests of engineering viability, and probes for reduced sensor deformation.
@@ 5.1 Ad hoc attachment of interposer to module
Figure 11 displays the early ad hoc procedure for attaching interposers onto flexes before assembling into a module. The SE-4445 is a two-part mixture that is hand mixed in a 1:1 ratio by weight. The Kapton film is oversized by 1 mm all-around to ensure full flex coverage while avoiding SE-4445 spilling onto the sensor surface. We initially use a mechanical hybrid and powerboard to test this procedure before using electrical counterparts, verifying the target glue thickness is achievable while ensuring satisfactory glue coverage. The Kapton film is manually cut to shape using scissors. Existing stencils (designed to deposit the F112 epoxy glue layer) are repurposed to deposit the SE-4445 onto the flexes using a steel ruler as a squeegee. After lifting the stencil, the operator deposits the Kapton film manually onto the SE-4445. An Essdee ink roller [66] is procured to apply pressure and ensure full coverage of SE-4445 by visual inspection. We then use existing tooling shimmed to account for the added m thickness for assembly onto a sensor. A pre-production ATLAS18SS sensor is used with good high-voltage characteristics.
@@ 5.2 Quality control of module with ad hoc interposer
Glue height metrology. Figure 10(b) shows the glue height metrology for SS-04. Circle markers show the target heights expected for the interposer stackup. This verifies our ad hoc stencilling procedure can control the deposition of the SE-4445, Kapton, and epoxy. Cross markers show the glue height measurements after MTC at two temperature ranges. The differences are negligible, confirming no undesirable delamination.
@@
(a) Current-voltage (IV) scans (b) Per-channel input noise Figure 13: Electrical test results for first interposer SS-04 module. (a) Current–voltage scans with (without) ‘AMAC’ occur before bonding up to 550 V (700 V). Pre-TC (warm IV) occurs just before thermocycling starts at C. Post-TC (cold IV) scans occur at the temperature after the noted run. (b) Every three-point gain input noise result during C thermocycling is shown for the hybrid X under stream. Oranges (blues) indicate tests above (below) C; lighter (darker) lines show earlier (later) testing. -
@@
The ratio of output noise and gain determines the input noise for each channel. Two tests at C occur before and after the main MTC sequence as baseline cross-checks. Pre-production SS modules typically exhibit anomalous excesses of input noise during C testing, which is the hallmark of cold noise [52]. This is not observed for SS-04. The severity of cold noise increases with (i) lower temperatures and (ii) doubling the power through the powerboard [52]. To test (i), we perform C testing and results for one set of readout channels. To test (ii), a “cold shunt test” automatically occurs immediately after the first and last cold test in the standard MTC sequence.
@@
(a) Before MTC (b) C (c) C (d) C Figure 14: Sensor height metrology of SS-04 before and after thermocycling through the temperatures indicated. -
@@
To continue in-house QA studies, we additionally construct one long-strip (LS-12) and another short-strip (SS-05) interposer module. For these two modules, we attach the silicone gel and Kapton film to flexes using module assembly tooling with shims to improve procedural reliability. Once assembled, these modules undergo extended QA thermocycling to probe their mechanical resilience. We proceed directly to C after the first ten thermocycles to expedite thermal stress. Due to epoxy squeeze-out over the sensor bias rail, SS-05 suffers early breakdown at V so undergoes MTC with powered LV electronics but disconnected HV.
@@ Acknowledgments
We are very grateful for the many helpful and thoughtful discussions with various members of the ITk strip module collaboration. We especially thank Will Fawcett for earlier collaboration on module pre-production, and Richard Shaw for engineering support in constructing the high-voltage supply and thermal interlock. The ATLAS Inner Tracker Strips Phase II upgrade at the Cavendish Laboratory in Cambridge is supported by the Science and Technology Facilities Council (STFC) of UK Research and Innovation (UKRI). The work of JL is generously supported by a Junior Research Fellowship at Trinity College, University of Cambridge. Copyright 2025 CERN for the benefit of the ATLAS ITk Collaboration. CC-BY-4.0 license.

### v2

@@ 1 Introduction
The successful upgrade of the ATLAS Experiment [1, 2] for the High-Luminosity Large Hadron Collider [3] underpins future discoveries in particle physics [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Its cornerstone is a new all-silicon Inner Tracker (ITk) detector [16, 17, 18, 19]. The ITk comprises pixel [17] and strip [18] sub-detectors, each designed as a barrel bookended by endcaps. The strips barrel requires detector units called modules, each comprising mm p-type silicon strip sensors fabricated by Hamamatsu Photonics K.K. [20]. Industrial adhesives [21, 22, 23] mechanically attach sensors to custom electronics supplying front-end power, readout, monitoring, and control [24, 25, 26, 27, 28, 29]. Module assembly and testing within operating expectations, called quality control (QC) [30, 31], are distributed worldwide.
An important production QC procedure is module thermal cycling (MTC). Modules must pass electrical tests during repeated cycles between two temperatures. The nominal operating window is between room (C) and end-of-life temperatures (C) to mitigate elevated leakage currents from radiation damage [32, 33, 34]. Every module undergoes ten thermocycles to verify electrical functionality and mechanical integrity between C. Cooling or power failures [35] can expose modules to temperatures outside the operating window. This motivates studies beyond nominal QC temperatures. Testing outside operating expectations is called quality assurance (QA).
However, we observe critical problems when thermocycling ten pre-production modules between C and C: visual inspection identifies a sensor fracture for one module (figure 1) and three modules develop excess leakage currents during C tests. Independent efforts report similar failure modes but only after loading modules onto local support structures, which is called the sensor fracturing problem [36, 37]. Simulation studies [38] suggest adverse stress from mismatched coefficients of thermal expansion (CTE) [39, 40, 41, 42] between the silicon sensor (fixed to local supports with adhesives) and copper in the flexible printed circuit board (flex PCB); the intermediate epoxy glue provides a high-modulus mechanical coupling that transfers stress.
@@ 2 Module assembly and setup
We briefly review the pre-production module assembly procedure and thermocycling setup. Figure 3(a) illustrates the original pre-production barrel module design. Pre-production assembly uses Loctite Eccobond F112 [23] epoxy resin111The ITk Strip community refers to this adhesive as TrueBlue due to its optical colour.. This attaches the sensor to hybrid(s) and powerboard PCBs housing custom microelectronics [54, 55, 56, 57] for front-end readout, control, and power. There are two types of modules: long strip (LS) that have one hybrid to read out 2560 channels, and short strip (SS) that have two hybrids to read out 5120 channels. After wire bonding, modules undergo QC testing [30] in the coldbox of the MTC setup situated at the University of Cambridge (figure 4). This apparatus combines thermoelectric Peltiers with liquid coolant circulated by a Grant chiller [58, 59, 60], interfaced with custom detector control system (DCS) and data acquisition (DAQ) [61, 62], with continuous environmental monitoring.
This paper probes mechanical deformations by measuring the sensor bow of free modules outside the MTC setup. We employ the Baty Venture 3030 coordinate-measuring machine (CMM) [63]. We develop an in-house script to convert the machine output into a Common Data Format (CDF) defined by standard QC procedures [30, 44]. A community-developed script then reads this CDF file and computes the sensor bow using a standard bow algorithm. The algorithm first performs a fit to all sensor points to define a sensor plane . The coordinate system convention is defined in the caption of figure 9. To remove any global sensor tilt, the residuals relative to the sensor plane are used to calculate the bow magnitude as the range of residuals:
@@ 3.2 High-voltage supply and interlock
We also develop a custom HV power supply as an add on to the MTC setup. It is based on a Spellman MPS DC-to-HV converter capable of delivering 2.5 mA at V. It uses an Arduino Leonardo microprocessor to operate the display, facilitate over-current and safety interlock functionality, and provide a serial-over-USB interface for remote readout and control. The HV converter has monitoring outputs for voltage and current, which are accurate to 1% of the full scale only. The unit features a dedicated interlock by interfacing to a magnetic reed switch attached to the coldbox lid (figure 5). When the lid is opened with HV enabled, the interlock rapidly reduces the HV output to 0 V at 100 V s-1. This introduces added safety beyond the existing software interlock.
@@ 3.3 Improved thermal chuck flatness
After fitting the shims and electrical connection pieces between the shims, feeler gauges determine the residual height differences with results in table 1. The post-shim values are all lower relative to the corresponding point before shimming, where the mean reduces to m with a maximum 170 m. This demonstrates improved chuck flatness relative to the original design. A bow of up to m is deemed tolerable for subsequent use. Comparative metrology measurements indicate a high degree of similarity between the bow of a module resting on a flat gauge plate and a module resting on a test frame bolted to a known flat jig surface. For consistency in the upcoming studies, sensor bow metrology is performed with the module resting on the same test frame at room temperature.
![[Uncaptioned image]](figs/chuck-abcd.jpg)
@@ 4.1 Test-to-destruction methodology
Extend thermocycling until fractures form. Sensor deformation from thermal stress accumulates in the sensor bow with repeated cycles, which can lead to sensor fractures. We evaluate electrical data and perform sensor visual inspection to search for fracture formation.
@@ 4.2 Bow evolution versus cycling temperature
We first report the regime with no sensor deformation. After the first cycle between C, we find a minor bow decrease of around m and the trends are correlated across all five modules, suggesting mechanical relaxation. For simplicity, we average the bow data over the five modules after C MTC to yield m (table 3). Thermocycling between C induces little change in bow, except LS-06 rises anomalously to m after the second C cycle due to accidental sensor damage. This damage comprises a mechanical chip in the sensor edge, where a small fragment including the bias rail around a few millimetres in width and length is physically detached; it is visible on the left edge of LS-06 as annotated in figure 9. Leaving the modules in dry storage for five days decreases the average bow by m, indicating mechanical relaxation. After C thermocycling, we find the mean bow reaches m, where the average of the individual increases is m relative to C cycling. The large standard deviation arises from the anomalous LS-06 value; excluding this outlier yields m. This establishes no statistically significant change in sensor deformation during MTC from C to C.
@@
This demonstrates that thermocycling above a critical temperature C “bakes in” sensor deformations permanently, enabling ex situ measurement outside the MTC setup. This hysteresis is also reported in endcap modules [44] and is hypothesised to arise from a glass transition of the epoxy adhesive [45] at C. While we lack the equipment for in situ metrology, off-site cold metrology of an unconstrained pre-production module shows sensor corners rises by about 1 mm at C, suggesting significant thermal stress during MTC. The mechanism of how the glass transition induces sensor deformation is not fully understood and remains an empirical observation for this paper.
@@ 4.3 Post-mortem of sensor fractures
Differences in fracture morphology. The LS-02 sensor distinctly fractures into multiple pieces, where we visually identify five major segments. We refer to the fractures that do not propagate from one edge to the opposite edge as secondary fractures. We observe that these develop perpendicular to the primary fracture left of the DC-DC converter shield box on the powerboard. One secondary fracture (medium blue) follows the powerboard glue edge with three near-right-angle kinks turning to the sensor top edge. The secondary fracture (dark blue) arises just above the primary fracture by the hybrid before turning perpendicularly at the top-left corner of the powerboard shield box to the sensor top edge, which does not trace a glue boundary.
@@ 5 Sensor deformation with stress-mitigating interposer
Simulation studies [38] suggest an order-of-magnitude reduction in stress by adding Kapton-silicone interposers. Due to its low modulus, the silicone gel reduces mechanical stress by decoupling the thermal deformation of the flexes from that of the sensor during cooling [47, 48]. Figure 10(a) schematically shows the flex-epoxy-sensor stackup augmented with 50 m of Kapton film and 100 m of SE-4445 silicone gel, whose concept is introduced in Refs. [38, 49]. The Kapton film is required to ensure sufficient adhesion given it is found that the epoxy does not adhere well to silicone. These two materials are already validated for use in the ITk detector because SE-4445 attaches modules onto local supports, whose electrical services supplied by bus tapes are made of Kapton [66]. Introducing these extra thin layers across all modules has a negligible impact on ITk material budget. This motivates us to prototype interposers in our laboratory by developing in-house designs, ad hoc assembly procedures, tests of engineering viability, and probes for reduced sensor deformation.
@@ 5.1 Ad hoc attachment of interposer to module
Figure 11 displays the early ad hoc procedure for attaching interposers onto flexes before assembling into a module. The SE-4445 is a two-part mixture that is hand mixed in a 1:1 ratio by weight. It is found that SE-4445 contamination on sensors induces undesirable early HV breakdown. The Kapton film is therefore oversized by 1 mm all-around to ensure full flex coverage while avoiding SE-4445 spilling onto the sensor surface. We initially use a mechanical hybrid and powerboard to test this procedure before using electrical counterparts, verifying the target glue thickness is achievable while ensuring satisfactory glue coverage. The Kapton film is manually cut to shape using scissors. Existing stencils (designed to deposit the F112 epoxy glue layer) are repurposed to deposit the SE-4445 onto the flexes using a steel ruler as a squeegee. After lifting the stencil, the operator deposits the Kapton film manually onto the SE-4445. An Essdee ink roller [67] is procured to apply pressure and ensure full coverage of SE-4445 by visual inspection. Upon SE-4445 compression, the fill factor of the repurposed stencil approximately achieves the target thickness that is later verified by metrology. We then use existing tooling shimmed to account for the added m thickness for assembly onto a sensor. A pre-production ATLAS18SS sensor is used with good high-voltage characteristics.
@@ 5.2 Quality control of module with ad hoc interposer
Glue height metrology. Figure 10(b) shows the glue height metrology for SS-04. Circle markers show the target heights expected for the interposer stackup. This verifies our ad hoc stencilling procedure can control the deposition of the SE-4445, Kapton, and epoxy. A small minority of points fall outside tolerance, but has negligible impact on downstream assembly such as wire bonding so is deemed unproblematic. Cross markers show the glue height measurements after MTC at two temperature ranges. The differences are negligible, confirming no undesirable delamination.
@@
(a) Current-voltage (IV) scans
(b) Per-channel input noise Figure 13: Electrical test results for first interposer SS-04 module. (a) Current–voltage scans with (without) ‘AMAC’ occur before bonding up to 550 V (700 V). Pre-TC (warm IV) occurs just before thermocycling starts at C. Post-TC (cold IV) scans occur at the temperature after the noted run. (b) Every three-point gain input noise result during C thermocycling is shown for the hybrid X under stream. Oranges (blues) indicate tests above (below) C; lighter (darker) lines show earlier (later) testing. -
@@
The ratio of output noise and gain determines the input noise for each channel. Two tests at C occur before and after the main MTC sequence as baseline cross-checks. Pre-production SS modules typically exhibit anomalous excesses of input noise during C testing, which is the hallmark of cold noise [53]. This is not observed for SS-04. The severity of cold noise increases with (i) lower temperatures and (ii) increasing the power load on the DC-DC converter on the powerboard as described in Ref. [53]. To test (i), we perform C testing and results for one set of readout channels. To test (ii), a “cold shunt test” occurs immediately after the first and last cold test in the standard MTC sequence, which temporarily shunts the current on the front-end chips to increase its power draw and thus load on the DC-DC converter.
@@
(a) Before MTC
(b) C
(c) C
(d) C Figure 14: Sensor height metrology of SS-04 before and after thermocycling through the temperatures indicated. -
@@
To continue in-house QA studies, we additionally construct one long-strip (LS-12) and another short-strip (SS-05) interposer module. For these two modules, we attach the silicone gel and Kapton film to flexes using module assembly tooling with shims to improve procedural reliability. The vacuum tooling is repurposed with shims to control consistency of the SE-4445 target height during curing. Once assembled, these modules undergo extended QA thermocycling to probe their mechanical resilience. We proceed directly to C after the first ten thermocycles to expedite thermal stress. Due to epoxy squeeze-out over the sensor bias rail, SS-05 suffers early breakdown at V so undergoes MTC with powered LV electronics but disconnected HV.
@@ Acknowledgments
We are very grateful for the many helpful and thoughtful discussions with various members of the ITk strip module collaboration. We especially thank Will Fawcett for earlier collaboration on module pre-production, and Richard Shaw for engineering support in constructing the high-voltage supply and thermal interlock. The ATLAS Inner Tracker Strip Phase II upgrade at the Cavendish Laboratory in Cambridge is supported by the Science and Technology Facilities Council (STFC) of UK Research and Innovation (UKRI). The work of JL is generously supported by a Junior Research Fellowship at Trinity College, University of Cambridge. Copyright 2025 CERN for the benefit of the ATLAS ITk Collaboration. CC-BY-4.0 license.